OpenStudy (anonymous):

A car travels at an average speed of 68 miles per hour. How long does it take to travel 306 miles?

jimthompson5910 (jim_thompson5910):

use the formula \[\Large D = r*t\]

jimthompson5910 (jim_thompson5910):

D = distance, which in this case is 306 miles r = rate (aka speed) which is 68 mph t = unknown but you use the formula above to solve for t

jimthompson5910 (jim_thompson5910):

D = r*t 306 = 68*t .... plug in the known values solve for t

OpenStudy (anonymous):

4.5

jimthompson5910 (jim_thompson5910):

yep so it will take 4.5 hours (4.5 hrs = 4 hrs, 30 minutes)